OpenStudy (anonymous):

two cars left a service center at the same time. One car travelled in one direction at 75 km/h. The other car travelled in the opposite direction at 85 km/h. After how long were they 600 km apart? can someone please explain how to do thesekind of questions and step by step please :)

OpenStudy (hoblos):

(75 +85)t =600 => t=600/160 = 3.75